The block weighing 8 kg started moving at an initial speed of 6 m / s along a rough horizontal surface. The intensity of the friction force is 12 N. Approximately determine the path that will traverse the block until the moment of complete stopping.


1
Expert's answer
2021-03-23T11:13:11-0400
"v^2=2as\\v^2=2\frac{F}{m}s\\6^2=2\frac{12}{8}s\\ s=12\ m"